# Don’t be lazy: CompleteP enables compute-efficient deep transformers

### v2

@@ Appendix D Scalings that Guarantee Our Desiderata
D.1 Desideratum 1: Stable Initialization
@@
The weights are initialized with $W^{\ell}_{ij}\sim\mathcal{N}(0,\frac{\sigma^{2}_{W}}{N})$. Defining $H^{\ell}\equiv\frac{1}{N}|\mathbf{h}^{\ell}|^{2}$ as the norm of the residual stream variables, we have the following recursion for $H^{\ell}$ at large $N$
@@
D.2 Desideratum 2: Maximal Update
@@
where $\eta^{\ell}$ is the learning rate for this hidden layer. Note that AdamW will contain the same scaling (with $N,L$) for the numerator and denominator but will also contain momentum (which does not change the scaling with $N,L$) and an additional $\epsilon$ parameter which we discuss how to scale with $N,L$ in Appendix E.4.
@@
Appendix E Derivation for bias, LN, WD, and AdamW $\epsilon$ parameterization adjustments
E.1 Bias Parameters
@@ E.3 Weight Decay
where the entries of $\mathbf{h}^{\ell+1}$ are $\Theta_{N,L}(1)$ by the arguments of Section D.1. Following the arguments of Section D.2, the learning rate $\eta$ must be set as
@@
to ensure the correct change to the residual stream variables. Since we desire the final term to be on the same scale this means that $\lambda=\Theta_{N,L}(1)$. Our proposed in weight decay scaling rule (Table 1) is also used in a few recent works [84, 98, 99]. The contribution of this section is to show no additional depth correction to the weight decay parameter $\lambda$ is required.
E.4 Adam $\epsilon$ Parameter
@@
Appendix F Depth-wise transfer of AdamW $\epsilon$
Everett et al. [38] show appropriate AdamW $\epsilon$ scaling is important for compute-efficient large models. Building on the width adjustment for AdamW $\epsilon$ introduced by Yang and Littwin [100], we add a depth correction for AdamW $\epsilon$. Figure 8 shows this correction enables stable training for a wider range of $\epsilon_{\text{base}}$.
@@ Appendix F Depth-wise transfer of AdamW $\epsilon$
Appendix G Depth Scaling Training Dynamics
We illustrate the training dynamics in the form of loss curves in Figure 9 and 10, which slice the $\eta$ transfer test in Figure 2 at CompleteP ($\alpha=1$)’s and $\alpha=0.5$’s optimal $\eta$ respectively. CompleteP ($\alpha=1$) consistently achieves lower training loss when model depth scales up. In $\alpha=0.5$, we point out the loss "crossing" behavior where deeper models continually take longer steps to scale better than shallower models. This aligns with our observation in Figure 2 where $\alpha=0.5$ desires larger $\eta$ with increasing model depth and laziness (defined in Definition).
Figure 9: Training dynamics of Figure 2 for all parameterizations at CompleteP ($\alpha=1$)’s optimal $\eta=2^{-8}$.
Figure 10: Training dynamics of Figure 2 for all parameterizations at $\eta=2^{-6}$, where $\alpha=0.5$ first demonstrates deeper-model-lower-loss ordering.
Appendix H 200 TPP parameterization comparison
Given the popularity of TPP>20 training for enhanced parameter efficiency [56], we also compare compare parameterizations at 200 TPP for $P_{\text{non-emb}}\in\{50M,300M\}$ in Table 3. CompleteP ($\alpha=1$) consistently achieves the best loss across all parameterizations, even in the 200TPP regime.
Table 3: Models trained for 200 tokens per parameters. Reporting SlimPJ validation cross-entropy loss. CompleteP $\alpha=1$ achieves better losses than µP and $\alpha=0.5$.
2.867 2.874 2.859 2.973 2.998 2.950 2.451 2.455 2.443 2.537 2.534 2.497
Appendix I Additional experimental details
Table 4 contains extensive details for all experiments. The rest of this section describes the remaining experimental details.
@@ Appendix I Additional experimental details
Figure $N$ $L$ $\sigma_{\text{init}}$ $\eta_{\text{base}}$ $\lambda_{\text{base}}$ $T_{\text{EMA}}$ $B$ Steps Tokens TPP
@@
3 256 Vary 0.02 Vary Vary 0.1407 Eqn. 47 Vary Vary 20
4, 13, 1 middle & right Vary Vary 0.02 2E-08 Vary 0.1407 Eqn. 47 Vary Vary 20
@@
11 Vary 6 0.09 Vary Vary Vary Eqn. 47 Vary Vary 20
12 Vary Vary 0.02 2E-08 Vary Vary Eqn. 47 Vary Vary 20
Table 3 Vary Vary 0.02 2E-08 Vary 0.1407 Eqn. 47 Vary Vary 200
I.1 General setup
For all experiments in this work, we train decoder-only Transformer language models [79] with pre-normalization, untied embeddings, ALiBi position embeddings [101] and $\text{ReLU}^{2}$ nonlinearity [102, 103], using the AdamW optimizer with decoupled weight decay [104] with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e-16$. The learning rate $\eta$ schedule follows a linear warmup of $n_{\text{warmup\_steps}}$ (Equation 46) and then a linear decay to zero [105].
$n_{\text{warmup\_steps}}=\min(\text{int}(0.1*n_{\text{steps}}),\text{int}(375e6/(B\cdot 2048))$ (46)
We pretrain our models using an autoregressive loss (i.e. the next token prediction objective) on the SlimPajama dataset [80] with a maximal sequence length of 2048 tokens using the GPT-2 tokenizer [79]. All models use $d_{\text{head}}=64$ for each attention head and feedforward dimension $4N$. For all parameterizations, we use $\mathbf{Q}^{\top}\mathbf{K}/N$ as proposed in [5] to account for correlation between $Q$ and $K$. We use $N_{\text{base}}=256,L_{\text{base}}=2$ and scale parameters according to Table 1.
All experiments were performed using Cerebras CS-3 systems.
I.2 Compute-efficient setup
In Figure 4, to study the compute-optimal setting, first our training setup must represent compute-optimal. Following [3], we train models for 20 tokens per parameter (TPP) to ensure a compute-optimal tradeoff of parameters and tokens. We also scale our batch size according to a power law in FLOPS. Since, McCandlish et al. [81], Kaplan et al. [2] showed the critical batch size scales with a power law in loss, and loss scales with a power law in FLOPS, the critical batch size scales with a power law in FLOPS [82, 83]. We follow Equation 47 based on empirical fits and rounded to the nearest multiple of 8, where $B,F$ are batch size and training FLOPS respectively.
@@ I.2 Compute-efficient setup
We adopt the $\sigma_{\text{base}}=0.02$ and the optimal $\eta_{\text{base}}$ for $L=L_{\text{base}}=2$ from Figure 2: $\eta_{\text{base}}=0.0039$.
To ensure optimal weight decay $\lambda$ as we scale model and dataset size, we follow Wang and Aitchison [84] by setting $\lambda_{\text{base}}$ to maintain the optimal $\tau_{\text{EMA}}=(\eta_{\text{base}}\lambda_{\text{base}}n_{\text{steps}})^{-1}$. A major shortcoming of $\tau_{\text{EMA}}$ is that it does not take the learning rate schedule into account, even though learning rates are decayed in practice. Despite this we observe reliable transfer of $\tau_{\text{EMA}}$ across model sizes, provided we use the same learning rate schedule. For SP and µP, Figure 11 validates that the claims of Wang and Aitchison [84] that the optimal $\tau_{\text{EMA}}$ remains stable across different learning rates and weight decay values. We then tune $\tau_{\text{EMA}}$ for SP, µP, $\alpha=0.5$, and CompleteP ($\alpha=1$). Figure 12 shows $\tau_{\text{EMA}}=0.1407$ close to optimal for all parameterizations and this adopted throughout this paper. When a non-zero weight decay is adopted, we use $\tau_{\text{ema}},\eta_{\text{base}},n_{\text{steps}}$ to decide the $\lambda_{\text{base}}$ used in each run.
@@
I.3 N:L study details
Here we include additional experimental details for Figure 4. We vary $N,L$ to approximately maintain $P_{\text{non-emb}}=12N^{2}L$. Table 5 contains all model shapes, token counts, and FLOPs used in Figure 4. Approximations for training FLOPs like $F=6ND$ do not account for embedding, attention, LN, nonlinearity or bias FLOPs. We use a very granular function to measure FLOPs. Although we maintain $P_{\text{non-emb}}$, the FLOPs are not equal mainly because of embedding and attention costs. We train $P_{\text{non-emb}}\in\{50M,300M,1.5B\}$ models, matching sizes from [2], but all at 20 TPP. [2] used a similar study design, but with the same number of tokens for all runs, meaning it didn’t follow the compute-optimal prescription of [3].
@@ I.3 N:L study details
For each $(F,X)$ point in Figure 4d, we calculate the “Loss increase vs. $\hat{X}_{\alpha=1}$” as $d=\frac{X-\hat{X}_{\alpha=1}}{X}$ and plot it in Figure 4a-c. To understand how the FLOP savings in Figure 4e are calculated, first recall the functional form of $\hat{X}(F)=(F/a)^{-}b$. Inverting yields $F(X)=aX^{-\frac{1}{b}}$. Due to the scale invariance of power laws, we can calculate:
[…]
Figure 13 shows how Figure 4f was created by fitting cubic functions to data $\{x=\log(N:L),y=d-d^{\text{optimal}}_{\alpha=1}\}$ and finding intersections with $y=1$. Cubic functions were used instead of quadratics to enable better fits to functions without an axis of symmetry.
Figure 13: Plotted loss increase from best point. Fit cubic functions to log-transformed data. Found intersections with $y=1$ for Figure 4f.
I.4 Downstream evaluation details
We evaluated our models on downstream tasks using Eleuther Harness Evaluation (EEH) [106]. The evaluation details are reported in Table 6.
@@ I.4 Downstream evaluation details
I.5 Figure 6 details
To visualize the notion of laziness in Equation 4:
@@ I.5 Figure 6 details
where $\bm{\theta}=\mathbf{W}_{(2)}^{\ell}$. We use $\eta_{0}=0.0001$ and train on MNIST [107]. We provide the code to reproduce this plot at https://github.com/EleutherAI/nanoGPT-mup/blob/completep/completep_examples/relative_distance_from_linearization.ipynb.

### v3

@@ Appendix C Scalings that Guarantee Our Desiderata
C.1 Desideratum 1: Stable Initialization
@@
The weights are initialized with $W^{\ell}_{ij}\sim\mathcal{N}(0,\frac{\sigma^{2}_{W}}{N})$. Defining $H^{\ell}\equiv\frac{1}{N}|\mathbf{h}^{\ell}|^{2}$ as the norm of the residual stream variables, we have the following recursion for $H^{\ell}$ at large NN
@@
C.2 Desideratum 2: Maximal Update
@@
where $\eta^{\ell}$ is the learning rate for this hidden layer. Note that AdamW will contain the same scaling (with $N,L$) for the numerator and denominator but will also contain momentum (which does not change the scaling with $N,L$) and an additional $\epsilon$ parameter which we discuss how to scale with $N,L$ in Appendix D.5.
@@
Appendix D Derivation for bias, LN, WD, and AdamW $\epsilon$ parameterization adjustments
D.1 On Generalizability of the Parameterization
In the later parts of this subsection, we will discuss derivations of parameterizations for several different components of the architecture via a heuristic method, which we find very accurately represents the scaling important quantities. However, this remains limited as there are other architecture components that are not contained by our manuscript, which may require further adjustments and derivations.
With this in mind, we can divide up the future generalizations into three categories:
  1. 1.
    No modifications required. This includes LayerNorm at any position (pre, post, QK-norm etc.), adding biases at different locations, LR schedule, and other well-normalized Adam-like algorithms (e.g. SignGD, Adagrad, Lion), or position embeddings (learned, RoPE, ALiBi, NoPE, etc.). The reason is that these do not change the scale of both the forward pass and hidden layer updates due to training, with respect to width and depth. LayerNorm, for example, maintains the $\Theta(1)$ scale of each neuron, and as long as the update maintains the $\Theta(1)$ scale, then no changes are required.
  2. 2.
    Slight modifications are required and known. A good example of this is SGD, where the learning rate scaling is derived in Table 1 of Bordelon et al. [8] (up to an equivalent ABC-reparameterization). Notice here that compared to Adam, the learning rate of SGD needs to be scaled up due to the prefactors in front of both the weight matrices and the residual branch, where as these extra factors are canceled out in Adam due to normalization.
  3. 3.
    More theoretical derivations required. This includes MoE, long context, batch size (and schedule), gradient clipping, LAMB, and momentum. Some of these are a relatively straight forward exercise to derive (e.g. LAMB and other optimization algorithms), others are less straight forward. In particular, any scaling that requires increasing the number of data points and training steps along with width and depth remains theoretically challenging. Our approach in this work computes the theoretical scaling for finitely many data points and training steps, and testing whether or not hyperparameter transfer empirically, which has been yielding very good results (see e.g. Figure˜3).
For the third category, we would effectively need to derive the scale of the updates and the downstream effects for each new architectural modification. A good example is consider the effect of bias weights in the next subsection. Here biases $\mathbf{b}^{\ell}$ are updated such that it contributes the same amount of changes to $\mathbf{h}^{\ell+1}$ as the usual weights $\mathbf{W}^{\ell}$. In order to do that, we need to choose the appropriate initialization and learning scale to satisfy this requirement. All other modifications follow the same mechanism, and the rest of this section serve as an example derivation for all future generalizations.
D.2 Bias Parameters
@@ D.4 Weight Decay
where the entries of $\mathbf{h}^{\ell+1}$ are $\Theta_{N,L}(1)$ by the arguments of Section˜C.1. Following the arguments of Section˜C.2, the learning rate $\eta$ must be set as
@@
to ensure the correct change to the residual stream variables. Since we desire the final term to be on the same scale this means that $\lambda=\Theta_{N,L}(1)$. Our proposed in weight decay scaling rule (Table˜1) is also used in a few recent works [89, 103, 104]. The contribution of this section is to show no additional depth correction to the weight decay parameter $\lambda$ is required.
D.5 Adam $\epsilon$ Parameter
@@
Appendix E Depth-wise transfer of AdamW $\epsilon$
Everett et al. [38] show appropriate AdamW $\epsilon$ scaling is important for compute-efficient large models. Building on the width adjustment for AdamW $\epsilon$ introduced by Yang and Littwin [105], we add a depth correction for AdamW $\epsilon$. Figure˜8 shows this correction enables stable training for a wider range of $\epsilon_{\text{base}}$.
@@ Appendix E Depth-wise transfer of AdamW $\epsilon$
Appendix F Depth Scaling Training Dynamics
We illustrate the training dynamics in the form of loss curves in Figure 9 and 10, which slice the $\eta$ transfer test in Figure˜2 at CompleteP ($\alpha=1$)’s and $\alpha=0.5$’s optimal $\eta$ respectively. CompleteP ($\alpha=1$) consistently achieves lower training loss when model depth scales up. In $\alpha=0.5$, we point out the loss "crossing" behavior where deeper models continually take longer steps to scale better than shallower models. This aligns with our observation in Figure˜2 where $\alpha=0.5$ desires larger $\eta$ with increasing model depth and laziness (defined in ˜Definition).
Figure 9: Training dynamics of Figure˜2 for all parameterizations at CompleteP ($\alpha=1$)’s optimal $\eta=2^{-8}$.
Figure 10: Training dynamics of Figure˜2 for all parameterizations at $\eta=2^{-6}$, where $\alpha=0.5$ first demonstrates deeper-model-lower-loss ordering.
Appendix G Additional experimental details
Table˜4 contains extensive details for all experiments. The rest of this section describes the remaining experimental details.
@@ Appendix G Additional experimental details
Figure NN $L$ $\sigma_{\text{init}}$ $\eta_{\text{base}}$ $\lambda_{\text{base}}$ $T_{\text{EMA}}$ $B$ Steps Tokens TPP
@@
3 256 Vary 0.02 Vary Vary 0.1407 Eqn. 46 Vary Vary 20
4, 13, 1 middle & right Vary Vary 0.02 2E-08 Vary 0.1407 Eqn. 46 Vary Vary 20
@@
11 Vary 6 0.09 Vary Vary Vary Eqn. 46 Vary Vary 20
12 Vary Vary 0.02 2E-08 Vary Vary Eqn. 46 Vary Vary 20
Table˜3 Vary Vary 0.02 2E-08 Vary 0.1407 Eqn. 46 Vary Vary 200
G.1 Compute-efficient setup
In Figure˜4, to study the compute-optimal setting, first our training setup must represent compute-optimal. Following [3], we train models for 20 tokens per parameter (TPP) to ensure a compute-optimal tradeoff of parameters and tokens. We also scale our batch size according to a power law in FLOPS. Since, McCandlish et al. [86], Kaplan et al. [2] showed the critical batch size scales with a power law in loss, and loss scales with a power law in FLOPS, the critical batch size scales with a power law in FLOPS [87, 88]. We follow Equation 46 based on empirical fits and rounded to the nearest multiple of 8, where $B,F$ are batch size and training FLOPS respectively.
@@ G.1 Compute-efficient setup
We adopt the $\sigma_{\text{base}}=0.02$ and the optimal $\eta_{\text{base}}$ for $L=L_{\text{base}}=2$ from Figure˜2: $\eta_{\text{base}}=0.0039$.
To ensure optimal weight decay $\lambda$ as we scale model and dataset size, we follow Wang and Aitchison [89] by setting $\lambda_{\text{base}}$ to maintain the optimal $\tau_{\text{EMA}}=(\eta_{\text{base}}\lambda_{\text{base}}n_{\text{steps}})^{-1}$. A major shortcoming of $\tau_{\text{EMA}}$ is that it does not take the learning rate schedule into account, even though learning rates are decayed in practice. Despite this we observe reliable transfer of $\tau_{\text{EMA}}$ across model sizes, provided we use the same learning rate schedule. For SP and µP, Figure˜11 validates that the claims of Wang and Aitchison [89] that the optimal $\tau_{\text{EMA}}$ remains stable across different learning rates and weight decay values. We then tune $\tau_{\text{EMA}}$ for SP, µP, $\alpha=0.5$, and CompleteP ($\alpha=1$). Figure˜12 shows $\tau_{\text{EMA}}=0.1407$ close to optimal for all parameterizations and this adopted throughout this paper. When a non-zero weight decay is adopted, we use $\tau_{\text{ema}},\eta_{\text{base}},n_{\text{steps}}$ to decide the $\lambda_{\text{base}}$ used in each run.
@@
G.2 N:L study details
Here we include additional experimental details for Figure˜4. We vary $N,L$ to approximately maintain $P_{\text{non-emb}}=12N^{2}L$. Table˜5 contains all model shapes, token counts, and FLOPs used in Figure˜4. Approximations for training FLOPs like $F=6ND$ do not account for embedding, attention, LN, nonlinearity or bias FLOPs. We use a very granular function to measure FLOPs. Although we maintain $P_{\text{non-emb}}$, the FLOPs are not equal mainly because of embedding and attention costs. We train $P_{\text{non-emb}}\in\{50M,300M,1.5B\}$ models, matching sizes from [2], but all at 20 TPP. [2] used a similar study design, but with the same number of tokens for all runs, meaning it didn’t follow the compute-optimal prescription of [3].
@@ G.2 N:L study details
For each $(F,X)$ point in Figure˜4d, we calculate the “Loss increase vs. $\hat{X}_{\alpha=1}$” as $d=\frac{X-\hat{X}_{\alpha=1}}{X}$ and plot it in Figure˜4a-c. To understand how the FLOP savings in Figure˜4e are calculated, first recall the functional form of $\hat{X}(F)=(F/a)^{-}b$. Inverting yields $F(X)=aX^{-\frac{1}{b}}$. Due to the scale invariance of power laws, we can calculate:
$\text{FLOP savings vs. }\mu\text{P}=1-(1-(d_{\mu P}-d_{\alpha=1}))^{-1/b}$ (47)
Figure˜13 shows how Figure˜4f was created by fitting cubic functions to data $\{x=\log(N:L),y=d-d^{\text{optimal}}_{\alpha=1}\}$ and finding intersections with $y=1$. Cubic functions were used instead of quadratics to enable better fits to functions without an axis of symmetry.
Figure 13: Plotted loss increase from best point. Fit cubic functions to log-transformed data. Found intersections with $y=1$ for Figure˜4f.
G.3 Downstream evaluation details
We evaluated our models on downstream tasks using Eleuther Harness Evaluation (EEH) [106]. The evaluation details are reported in Table˜6.
@@ G.3 Downstream evaluation details
G.4 Figure 6 details
To visualize the notion of laziness in Equation˜4:
@@ G.4 Figure 6 details
where $\bm{\theta}=\mathbf{W}_{(2)}^{\ell}$. We use $\eta_{0}=0.0001$ and train on MNIST [107]. We provide the code to reproduce this plot in supplementary material.